# Residual lifetime prediction for heterogeneous degradation data by Bayesian semi-parametric method

### v1

@@ 3 Residual lifetime distribution
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P[\alpha^{(i)}+\beta_{new}^{(i)}(t+t_{k})+\epsilon_{new,t+t_{k}}\geq D|\alpha=\alpha^{(i)},\beta_{new}=\beta_{new}^{(i)},\sigma_{\epsilon}^{2}=(\sigma_{\epsilon}^{2})^{(i)}]$
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P[\epsilon_{new,t+t_{k}}\geq D-\alpha^{(i)}-\beta_{new}^{(i)}(t+t_{k})|\alpha=\alpha^{(i)},\beta_{new}=\beta_{new}^{(i)},\sigma_{\epsilon}^{2}=(\sigma_{\epsilon}^{2})^{(i)}]$
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P[\epsilon_{{new,t+t_{k}}}\geq D-\alpha^{(i)}-\beta_{new}^{(i)}(t+t_{k})|\sigma_{\epsilon}^{2}=(\sigma_{\epsilon}^{2})^{(i)}]\hskip 5.69054pt$
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P\Big{[}\frac{\epsilon_{new,t+t_{k}}}{\sigma_{\epsilon}^{(i)}}\geq\frac{D-\alpha^{(i)}-\beta_{new}^{(i)}(t+t_{k})}{\sigma_{\epsilon}^{(i)}}|\sigma_{\epsilon}^{2}=(\sigma_{\epsilon}^{2})^{(i)}\Big{]}$
[…]
where $\Phi(\cdot)$ is cdf of standard normal random variable. So we approximate the residual lifetime distribution by the expression derived in equation (11). For the approximated distribution we have,
$F_{T_{new}|\bm{Y}=\bm{y}}^{approx}(0)=$ $P[T_{new}\leq 0|\bm{Y}=\bm{y}]$ (12)
@@
So, equation (12) implies that the random variable representing the residual lifetime of a new unit given the entire degradation data can take negative values. But we assumed that the support of $T_{new}$ is $(0,\infty)$. Because of this, we consider the distribution function for $T_{new}$ given $\bm{Y}=\bm{y}$ with an additional constraint $T_{new}>0$ which we denote as $F_{T_{new}|\bm{Y}=\bm{y},T_{new}>0}(t)$ and $F_{T_{new}|\bm{Y}=\bm{y},T_{new}>0}^{approx}(t)$ is the corresponding approximated distribution function. This expression of this distribution function is given as below:
$F_{T_{new}|\bm{Y}=\bm{y},T_{new}>0}(t)=$ $P[T_{new}\leq t|T_{new}>0,\bm{Y}=\bm{y}]$ (13)
$=$ $\frac{P{(T_{new}\leq t,T_{new}>0|\bm{Y}=\bm{y})}}{P{(T_{new}>0|\bm{Y}=\bm{y})}}$
$=$ $\frac{P{(0<T_{new}\leq t|\bm{Y}=\bm{y})}}{P{(T_{new}>0|\bm{Y}=\bm{y})}}$
$=$ $\frac{P{(T_{new}\leq t|\bm{Y}=\bm{y})-P(T_{new}<0|\bm{Y}=\bm{y})}}{1-P{(T_{new}\leq 0|\bm{Y}=\bm{y})}}$
$F_{T_{new}|\bm{Y}=\bm{y},T_{new}>0}^{approx}(t)=$ $\frac{\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\Phi\Big{[}\frac{\alpha^{(i)}+\beta_{new}^{(i)}(t+t_{k})-D}{\sigma_{\epsilon}^{(i)}}\Big{]}-\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\Phi\Big{[}\frac{\alpha^{(i)}+\beta_{new}^{(i)}t_{k}-D}{\sigma_{\epsilon}^{(i)}}\Big{]}}{1-\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\Phi\Big{[}\frac{\alpha^{(i)}+\beta_{new}^{(i)}t_{k}-D}{\sigma_{\epsilon}^{(i)}}\Big{]}}.$
The approximated pdf of $T_{new}|\bm{Y}=\bm{y},T_{new}>0$, is given by :
$f_{T_{new}|\bm{Y}=\bm{y},T_{new}>0}^{approx}(t)=\frac{\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\phi\Big{[}\frac{\alpha^{(i)}+\beta_{new}^{(i)}(t+t_{k})-D}{\sigma_{\epsilon}^{(i)}}\Big{]}\times\frac{\beta_{new}^{(i)}}{\sigma_{\epsilon}^{(i)}}}{1-\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\Phi\Big{[}\frac{\alpha^{(i)}+\beta_{new}^{(i)}t_{k}-D}{\sigma_{\epsilon}^{(i)}}\Big{]}}.$ (14)
where $\phi(\cdot)$is the pdf of standard normal random variable.
@@
Observe that if $\beta_{new}^{(i)}>0$ $\forall$ $i$, then $F_{T_{new}|\bm{Y}=\bm{y}}^{approx}(t)$, $F_{T_{new}|\bm{Y}=\bm{y},T_{new}>0}^{approx}(t)$ is a cumulative distribution function and $f_{T_{new}|\bm{Y}=\bm{y},T_{new}>0}^{approx}(t)$ is a probability density function. Observe that the approximated pdf of residual lifetime depends on the samples from the posterior distribution of $\alpha$, $\beta_{new}$ and $\sigma_{\epsilon}^{2}$. In this procedure entire observed degradation measures are utilized to derive residual lifetime distribution of a new unit. The proposed method has an advantage in the sense that is as soon as new degradation signals are available, we update the posterior distribution required parameters and sample from the updated posterior distribution and use the samples to estimate residual lifetime distribution. Next we consider how to generate samples from posterior distribution of required parameters using the proposed Bayesian semi-parametric degradation model.
@@ 4 Generating observations from posterior distribution
Here we consider simulation from the posterior distribution. We assume that $\alpha\sim\mathcal{N}(\mu_{\alpha},\sigma_{\alpha}^{2})$ and $\mu_{\alpha}\sim\mathcal{N}(\mu_{1},\sigma_{1}^{2})$, $(\sigma_{\alpha}^{2})^{-1}\sim\mathcal{G}(a_{1},a_{2})$, where we choose $(\mu_{1},\sigma_{1}^{2})=(0,10^{3})$ and $(a_{1},a_{2})=(0.001,0.001)$. It is considered that $\mu_{h}\sim\mathcal{N}(m_{\mu},\sigma_{z})$, $({\sigma_{z}^{2}})^{-1}\sim\mathcal{G}(\tau_{1},\tau_{2})$ , $({\sigma_{h}^{2}})^{-1}\sim\mathcal{G}(a,b)$, $h=1,.....,N$. Available prior information about the degradation rate can be used for choosing hyperparameter $m_{\mu}$. Santos and Loschi [26] used observed data for choosing the values of hyperparameters of $\mu_{h}$ and $\sigma_{h}^{2}$. Motivated by their ideas, we choose $m_{\mu}$, $a$ and $b$ as follows. We fit a simple linear regression model using the expression of equation (2) for the degradation path of each unit. Assuming that there are enough observations to estimate $\beta_{i}$ for each $i$, let $\beta_{i}^{*}$ be the its least square estimate of for $i$th unit, $i=1,...,(n+1)$. Now we consider $m_{\mu}$ and $v_{\beta}$ as the sample average and variance of the $(n+1)$ many $\beta_{i}^{*}$. For the hyper-parameters of prior distribution of $\sigma_{h}^{2}$, we consider $a=\sqrt{v_{\beta}},b={\sqrt[3]{v_{\beta}}}$. For the hyperpriors for $({\sigma_{z}^{2}})^{-1}$, we choose $(\tau_{1},\tau_{2})=(0.001,0.001)$ and for $({\sigma_{\epsilon}^{2}})^{-1}$, we consider $(a_{\epsilon},b_{\epsilon})=(0.001,0.001)$. For the Dirichlet mass parameter $\gamma$, it is assumed that $\gamma\sim G(\eta_{1},\eta_{2})$, where $\eta_{1}=2,\eta_{2}=2$, since this is a good choice as mentioned in Ishwaran [31]. Furthermore it is also assumed that prior distributions are independent.
@@ 4.1 Conditional distributions for Gibbs sampling
We generate observations from posterior distribution by Gibbs sampling method. Due to Dirichlet process prior there are ties among the classification variables. Let us consider that ${{K_{1}}^{*},.....,{K_{m}}^{*}}$ denotes the present $m\leq(n+1)$ unique values of K, where $\textbf{K}=(K_{1},.....,K_{n+1})$. In this article, we denote a vector $\textbf{Q}_{(-i)}$ which contains all $Q_{j}$’s where $j\neq i$. In each iteration of the Gibbs sampler we simulate from the following conditional distributions :
@@
    Conditional distribution of $\alpha$ given $(\bm{\beta},\mu_{\alpha},\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is
@@
    Conditional distribution of $\mu_{\alpha}$ given $(\bm{\beta},\alpha,\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is
    $\mathcal{N}\Big{(}\frac{\alpha\sigma_{1}^{2}+\mu_{1}\sigma_{\alpha}^{2}}{\sigma_{1}^{2}+\sigma_{\alpha}^{2}},\frac{\sigma_{\alpha}^{2}\sigma_{1}^{2}}{\sigma_{\alpha}^{2}+\sigma_{1}^{2}}\Big{)}$.
@@
    Conditional distribution of $(\sigma_{\alpha}^{2})^{-1}$ given $(\bm{\beta},\alpha,\mu_{\alpha},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is
    $\mathcal{G}\Big{(}a_{1}+\frac{1}{2},b_{1}+\frac{(\alpha-\mu_{\alpha})^{2}}{2}\Big{)}$.
@@
    Conditional distribution of $\beta_{i}$ given $(\bm{\beta}_{(-i)},\alpha,\mu_{\alpha},\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is
    $\mathcal{N}\Big{(}\frac{\sigma_{K_{i}}^{2}\sum\limits_{j=1}^{n_{i}}t_{ij}(y_{ij}-\alpha)+\sigma_{\epsilon}^{2}\mu_{K_{i}}}{\sigma_{K_{i}}^{2}\sum\limits_{j=1}^{n_{i}}t_{ij}^{2}+\sigma_{\epsilon}^{2}},\frac{\sigma_{K_{i}}^{2}\sigma_{\epsilon}^{2}}{\sigma_{K_{i}}^{2}\sum\limits_{j=1}^{n_{i}}t_{ij}^{2}+\sigma_{\epsilon}^{2}}\Big{)},\hskip 14.22636pti=1,...,n+1$.
@@
    Conditional distribution of $(\sigma_{\epsilon}^{2})^{-1}$ given $(\bm{\beta},\alpha,\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\gamma,\bm{Y})$ is,
@@
    Conditional distribution of $K_{i}$ given $(\bm{\beta},\alpha,\sigma_{\alpha}^{2},\bm{\mu},\sigma_{z}^{2},\textbf{K}_{(-i)},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is,
    $\sum\limits_{h=1}^{N}p_{h,i}\delta_{k}(\cdot)$,       $i=1,...,n+1$,
    where $(p_{1,i},......,p_{N,i})\propto(\frac{p_{1}}{\sigma_{1}}\exp(-\frac{(\beta_{i}-\mu_{1})^{2}}{2\sigma_{1}^{2}}),.....,\frac{p_{N}}{\sigma_{N}}\exp(-\frac{(\beta_{i}-\mu_{N})^{2}}{2\sigma_{N}^{2}}))$.

### v2

@@ 3 Residual lifetime distribution
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P[\alpha^{(i)}+\beta_{new}^{(i)}(t+t_{k})+\epsilon_{new}\geq D|\alpha=\alpha^{(i)},\beta_{new}=\beta_{new}^{(i)},\sigma_{\epsilon}^{2}=(\sigma_{\epsilon}^{2})^{(i)}]$
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P[\epsilon_{new}\geq D-\alpha^{(i)}-\beta_{new}^{(i)}(t+t_{k})|\alpha=\alpha^{(i)},\beta_{new}=\beta_{new}^{(i)},\sigma_{\epsilon}^{2}=(\sigma_{\epsilon}^{2})^{(i)}]$
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}P[\epsilon_{{new}}\geq D-\alpha^{(i)}-\beta_{new}^{(i)}(t+t_{k})|\sigma_{\epsilon}^{2}=(\sigma_{\epsilon}^{2})^{(i)}]\hskip 5.69054pt$
$=$ $\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\Phi\Big[\frac{\alpha^{(i)}+\beta_{new}^{(i)}(t+t_{k})-D}{\sigma_{\epsilon}^{(i)}}\Big],\hskip 145.6781pt(12)$
$f_{T_{new}|\bm{Y}}^{approx}(t)={\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\phi\Big[\frac{\alpha^{(i)}+\beta_{new}^{(i)}(t+t_{k})-D}{\sigma_{\epsilon}^{(i)}}\Big]\times\frac{\beta_{new}^{(i)}}{\sigma_{\epsilon}^{(i)}}}.$ (13)
where $\Phi(\cdot)$, $\phi(\cdot)$ are respectively cdf, pdf of standard normal random variable. So we approximate the residual lifetime distribution by the expression derived in equation (12). For the approximated distribution we have,
$F_{T_{new}|\bm{Y}}^{approx}(0)=$ $P[T_{new}\leq 0|\bm{Y}=\bm{y}]$ (14)
@@
Equation (14) implies that the random variable representing the residual lifetime of a new unit given the entire degradation data can take negative values with respect to the approximated distribution. Similarly, the same phenomena follows for the residual lifetime distribution given by equation (10). But we assumed that the support of $T_{new}$ is $(0,\infty)$. Because of this, we consider to derive the residual lifetime distribution function for the new unit with an additional constraint $T_{new}>0$ which we denote as $F_{T_{new}|\bm{Y},T_{new}>0}(t)$ given $\bm{Y}=\bm{y}$ with an additional constraint $T_{new}>0$ which we denote as $F_{T_{new}|\bm{Y},T_{new}>0}(t)$ and $F_{T_{new}|\bm{Y}=\bm{y},T_{new}>0}^{approx}(t)$ is the corresponding approximated distribution function. This expression of this distribution function is given as below:
$F_{T_{new}|T_{new}>0,\alpha,\beta_{new},\sigma_{\epsilon}^{2}}(t)=$ $P[T_{new}\leq t|T_{new}>0,\alpha=\alpha^{*},\beta_{new}=\beta_{new}^{*},\sigma_{\epsilon}^{2}=\sigma_{\epsilon}^{2*}]$ (15)
$=$ $\frac{P{(0<T_{new}\leq t|\alpha=\alpha^{*},\beta_{new}=\beta_{new}^{*},\sigma_{\epsilon}^{2}=\sigma_{\epsilon}^{2*})}}{1-P{(T_{new}\leq 0|\alpha=\alpha^{*},\beta_{new}=\beta_{new}^{*},\sigma_{\epsilon}^{2}=\sigma_{\epsilon}^{2*})}}.$
$=$ $\frac{\Phi\Big[\frac{\alpha^{*}+\beta_{new}^{*}(t+t_{k})-D}{\sigma_{\epsilon}^{*}}\Big]-\Phi\Big[\frac{\alpha^{*}+\beta_{new}^{*}t_{k}-D}{\sigma_{\epsilon}^{*}}\Big]}{1-\Phi\Big[\frac{\alpha^{*}+\beta_{new}^{*}t_{k}-D}{\sigma_{\epsilon}^{*}}\Big]}$
$F_{T_{new}|\bm{Y},T_{new}>0}^{approx}(t)=$ $\frac{\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\Phi\Big[\frac{\alpha^{(i)}+\beta_{new}^{(i)}(t+t_{k})-D}{\sigma_{\epsilon}^{(i)}}\Big]-\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\Phi\Big[\frac{\alpha^{(i)}+\beta_{new}^{(i)}t_{k}-D}{\sigma_{\epsilon}^{(i)}}\Big]}{1-\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\Phi\Big[\frac{\alpha^{(i)}+\beta_{new}^{(i)}t_{k}-D}{\sigma_{\epsilon}^{(i)}}\Big]}.$ (16)
The conditional pdf of $T_{new}$ given $\bm{Y}=\bm{y}$ and $T_{new}>0$ with respect to the approximated distribution function given by equation (16), is:
$f_{T_{new}|\bm{Y},T_{new}>0}^{approx}(t)=\frac{\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\phi\Big[\frac{\alpha^{(i)}+\beta_{new}^{(i)}(t+t_{k})-D}{\sigma_{\epsilon}^{(i)}}\Big]\times\frac{\beta_{new}^{(i)}}{\sigma_{\epsilon}^{(i)}}}{1-\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\Phi\Big[\frac{\alpha^{(i)}+\beta_{new}^{(i)}t_{k}-D}{\sigma_{\epsilon}^{(i)}}\Big]}.$ (17)
@@
Observe that if $\beta_{new}^{(i)}>0$ $\forall$ $i$, then $F_{T_{new}|\bm{Y}}^{approx}(t)$, $F_{T_{new}|\bm{Y},T_{new}>0}^{approx}(t)$, given by equations (12) and (16) is a cumulative distribution function and $f_{T_{new}|\bm{Y},T_{new}>0}^{approx}(t)$, $f_{T_{new}|\bm{Y}}^{approx}(t)$ given by equation (13) and (17) is a probability density function. Similarly, the expression derived in equation (15) is a distribution function when $\beta_{new}^{*}>0$. Kolmogorov-Smirnov (KS) distance can be used to measure the closeness of the approximated distribution and the true distribution. Observe that the approximated pdf of residual lifetime depends on the samples from the posterior distribution of $\alpha$, $\beta_{new}$ and $\sigma_{\epsilon}^{2}$. In this procedure entire observed degradation measures are utilized to derive residual lifetime distribution of a new unit. The proposed method has an advantage in the sense that, as soon as new degradation signals are available, we update the posterior distribution of required parameters and sample from the updated posterior distribution and use the samples to estimate residual lifetime distribution. Next we consider how to generate samples from posterior distribution of required parameters using the proposed Bayesian semi-parametric degradation model.
@@ 4 Generating observations from posterior distribution
Here we consider simulations from the posterior distribution. To check the impact of prior distribution on the posterior estimates for the proposed semi-parametric model, we consider 5 different prior combinations for parameters, which are given in Table 1. Similarly, we also consider different prior combinations for the parametric model which are presented in Table 2.
Table 1: Prior Specifications for Bayesian semi-parametric model
Scenario              Prior Distribution
1 $\alpha\sim\mathcal{N}(0,1000)$, $\mu_{h}\sim\mathcal{N}(0,\sigma_{z}^{2})$, $({\sigma_{h}^{2}})^{-1}\sim\mathcal{G}(1,0.01)$, $h=1,\dots,N$, $({\sigma_{z}^{2}})^{-1}\sim\mathcal{G}(0.01,0.01)$, $({\sigma_{\epsilon}^{2}})^{-1}\sim\mathcal{G}(0.01,0.01)$, $\gamma\sim\mathcal{G}(0.01,0.01)$.
2 $\alpha\sim\mathcal{N}(0,1000)$, $\mu_{h}\sim\mathcal{N}(m_{\mu},\sigma_{z}^{2})$, $({\sigma_{h}^{2}})^{-1}\sim\mathcal{G}(a,b)$, $h=1,\dots,N$, $({\sigma_{z}^{2}})^{-1}\sim\mathcal{G}(0.01,0.01),({\sigma_{\epsilon}^{2}})^{-1}\sim\mathcal{G}(0.01,0.01)$, $\gamma\sim\mathcal{G}(2,2)$.
3 $\alpha\sim\mathcal{N}(0,1000)$, $\mu_{h}\sim\mathcal{N}(m_{\mu},\sigma_{z}^{2})$, $({\sigma_{h}^{2}})^{-1}\sim\mathcal{G}(a,b)$, $h=1,\dots,N$, $({\sigma_{z}^{2}})^{-1}\sim\mathcal{G}(0.01,0.01),({\sigma_{\epsilon}^{2}})^{-1}\sim\mathcal{G}(0.01,0.01)$, $\gamma\sim\mathcal{G}(0.01,0.01)$.
4 $\alpha\sim\mathcal{N}(0,1000)$, $\mu_{h}\sim\mathcal{N}(m_{\mu},\sigma_{z}^{2})$, $\sigma_{h}\sim\text{Half-Cauchy}(A_{1})$, $h=1,\dots,N$, ${\sigma_{z}}\sim\text{Half-Cauchy}(A_{2}),({\sigma_{\epsilon}^{2}})^{-1}\sim\mathcal{G}(0.01,0.01)$, $\gamma\sim\mathcal{G}(2,2)$.
5 $\alpha\sim\mathcal{N}(0,1000)$, $\mu_{h}\sim\mathcal{N}(m_{\mu},\sigma_{z}^{2})$, $\sigma_{h}\sim\text{Half-Cauchy}(A_{1})$, $h=1,\dots,N$, $\sigma_{z}\sim\text{Half-Cauchy}(A_{2}),({\sigma_{\epsilon}^{2}})^{-1}\sim\mathcal{G}(0.01,0.01)$, $\gamma\sim\mathcal{G}(0.01,0.01)$.
Table 2: Prior Specifications for Bayesian parametric model
Scenario              Prior Distribution
1 $\alpha\sim\mathcal{N}(0,1000)$, $\mu_{\beta}\sim\mathcal{N}(0,\sigma_{z}^{2})$, $({\sigma_{\beta}^{2}})^{-1}\sim\mathcal{G}(0.01,0.01)$, $({\sigma_{z}^{2}})^{-1}\sim\mathcal{G}(0.01,0.01),({\sigma_{\epsilon}^{2}})^{-1}\sim\mathcal{G}(0.01,0.01)$.
2 $\alpha\sim\mathcal{N}(0,1000)$, $\mu_{\beta}\sim\mathcal{N}(m_{\mu},\sigma_{z})$, $({\sigma_{z}^{2}})^{-1}\sim\mathcal{G}(0.01,0.01)$, $({\sigma_{\beta}^{2}})^{-1}\sim\mathcal{G}(a,b),({\sigma_{\epsilon}^{2}})^{-1}\sim\mathcal{G}(0.01,0.01)$.
3 $\alpha\sim\mathcal{N}(0,1000)$, $\mu_{\beta}\sim\mathcal{N}(m_{\mu},\sigma_{z}^{2})$, $\sigma_{\beta}\sim\text{Half-Cauchy}(A_{1})$, $\sigma_{z}\sim\text{Half-Cauchy}(A_{2}),({\sigma_{\epsilon}^{2}})^{-1}\sim\mathcal{G}(0.01,0.01)$.
@@
For the semi-parametric model, we consider vague priors for fixed effect $\alpha$, inverse of error variance $\sigma_{\epsilon}^{2}$ and concentration parameter $\gamma$ of Dirichlet process in the first scenario. In scenario 2, we consider informative priors for mean $\mu_{h}$ and inverse of variance $(\sigma_{h}^{2})^{-1}$. Santos and Loschi [39] used observed data for choosing the values of hyperparameters for prior distributions of $\mu_{h}$ and $(\sigma_{h}^{2})^{-1}$. Motivated by their ideas, we choose $m_{\mu}$, $a$ and $b$ as follows. We fit a simple linear regression model assuming $\eta(t_{ij};\bm{\alpha},\bm{\beta}_{i})=\alpha+\beta_{i}t_{ij}$, for the degradation path of each unit. Assuming that there are enough observations to estimate $\beta_{i}$ for each $i$, let $\beta_{i}^{*}$ be its least square estimate of for $i$th unit, $i=1,\dots,(n+1)$. Now we consider $m_{\mu}$ and $v_{\beta}$ as the sample average and variance of the $(n+1)$ many $\beta_{i}^{*}$. For the hyper-parameters of prior distribution of $(\sigma_{h}^{2})^{-1}$, we consider $a=\sqrt{v_{\beta}},b={\sqrt[3]{v_{\beta}}}$. Ishwaran and James [49] used a informative prior for $\gamma$ for normal mean mixture model. In this article, we consider informative prior $\mathcal{G}(2,2)$ for $\gamma$ for prior 2, whereas vague prior $\mathcal{G}(0.01,0.01)$ is considered for $\gamma$ in scenario 3. A vague gamma prior is considered for $({\sigma_{z}^{2}})^{-1}$ for the first three scenarios, where the hyperpriors are taken as $(0.01,0.01)$. In a hierarchical model, Gelman ([40]) introduced half-Cauchy prior for the standard deviation parameter which is denoted by Half-Cauchy(A). We consider half-cauchy distribution for the standard deviation $(\sigma_{h})^{-1}$ and $(\sigma_{z})^{-1}$ with $A_{1}=A_{2}=25$. Prior 4 and 5 is constructed with these half-cauchy distribution for $(\sigma_{h})^{-1}$ and $(\sigma_{z})^{-1}$ along with consideration of informative and vague prior for the concentration parameter $\gamma$ respectively. For each scenario we consider $({\sigma_{\epsilon}^{2}})^{-1}\sim\mathcal{G}(0.01,0.01)$.
For the Bayesian parametric model, we consider three different prior scenarios. For Bayesian parametric method, we consider $\beta_{i}\sim N(\mu_{\beta},\sigma_{\beta}^{2})$, and $\mu_{\beta}\sim\mathcal{N}(0,\sigma_{z}^{2})$ and $(\sigma_{\epsilon}^{2})^{-1}\sim G(a_{\epsilon},b_{\epsilon})$. Firstly, we consider a non-informative prior for $\sigma_{\beta}^{2}$ and $\sigma_{z}^{2}$. For the second scenario we take informative prior for both mean and variance parameter of random effects. The hyperpriors $m_{\mu},a$ and b are chosen according to the procedure described by prior 2 for semi-parametric model. For prior 3 we consider half-Cauchy prior for $(\sigma_{h})^{-1}$ and $(\sigma_{z})^{-1}$ with $A_{1}=A_{2}=25$. Similar to the semi-parametric method, we consider a vague prior for $({\sigma_{\epsilon}^{2}})^{-1}\sim\mathcal{G}(a_{\epsilon},b_{\epsilon})$, with $(a_{\epsilon},b_{\epsilon})=(0.001,0.001)$ for each scenario. Furthermore, it is also assumed that parameters are independent with respect to prior distributions for both parametric and semi-parametric model.
@@ 4.1 Conditional distributions for Gibbs sampling
To simulate observations from the posterior distribution given by (7), we use the hierarchical structure of the model described by the expression in(6). To generate observations from posterior distribution, we use Gibbs sampling method. It may be noted that due to Dirichlet process prior there are ties among the classification variables. Let us consider that ${K_{1}}^{*},\dots,{K_{m}}^{*}$} denotes $m\leq(n+1)$ unique values of K, where $\textbf{K}=(K_{1},\dots,K_{n+1})$. In this article, we denote a vector $\textbf{Q}_{(-i)}$ which contains all $Q_{j}$’s where $j\neq i$. In each iteration of the Gibbs sampler we simulate from the following conditional distributions :
@@
    Conditional distribution of $\alpha$ given $(\bm{\beta},\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is
@@
    Conditional distribution of $\beta_{i}$ given $(\bm{\beta}_{(-i)},\alpha,\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is
    $\mathcal{N}\Big(\frac{\sigma_{K_{i}}^{2}\sum\limits_{j=1}^{n_{i}}t_{ij}(y_{ij}-\alpha)+\sigma_{\epsilon}^{2}\mu_{K_{i}}}{\sigma_{K_{i}}^{2}\sum\limits_{j=1}^{n_{i}}t_{ij}^{2}+\sigma_{\epsilon}^{2}},\frac{\sigma_{K_{i}}^{2}\sigma_{\epsilon}^{2}}{\sigma_{K_{i}}^{2}\sum\limits_{j=1}^{n_{i}}t_{ij}^{2}+\sigma_{\epsilon}^{2}}\Big),\hskip 14.22636pti=1,\dots,n+1$.
@@
    Conditional distribution of $(\sigma_{\epsilon}^{2})^{-1}$ given $(\bm{\beta},\alpha,\bm{\mu},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\gamma,\bm{Y})$ is,
@@
    Conditional distribution of $K_{i}$ given $(\bm{\beta},\alpha,\bm{\mu},\sigma_{z}^{2},\textbf{K}_{(-i)},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is,
    $\sum\limits_{h=1}^{N}p_{h,i}\delta_{k}(\cdot)$,     $i=1,\dots,n+1$,
    where $(p_{1,i},\dots,p_{N,i})\propto(\frac{p_{1}}{\sigma_{1}}\exp(-\frac{(\beta_{i}-\mu_{1})^{2}}{2\sigma_{1}^{2}}),\dots,\frac{p_{N}}{\sigma_{N}}\exp(-\frac{(\beta_{i}-\mu_{N})^{2}}{2\sigma_{N}^{2}}))$.
@@
    Conditional distribution of p given $(\bm{\beta},\bm{\mu},\sigma_{z}^{2},\textbf{K},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y})$ is represented in terms of distribution of $V_{h}^{*}$’s, $h=1,\dots,(N-1)$, where $p_{1}=V_{1}^{*}$ and $p_{h}=(1-V_{1}^{*})(1-V_{2}^{*})\dots(1-V_{h-1}^{*})V_{h}^{*}\hskip 2.84526pt,h=2,\dots,(N-1)$ and $p_{N}=1-\sum\limits_{h=1}^{N-1}p_{h}$ and the $V_{h}^{*}$ are independent and $V_{h}^{*}\sim Beta(1+r_{h},\gamma+\sum\limits_{l=k+1}^{N}r_{l})$, where $r_{h}=\sum\limits_{i=1}^{n}I_{K_{i}=h}$, $h=1,\dots,N$.
@@
    Conditional distribution of ${\mu}_{h}$ given $(\bm{\beta},\alpha,\bm{\mu}_{(-h)},\sigma_{z}^{2},\textbf{K},\textbf{p},\bm{\sigma}^{2},\sigma_{\epsilon}^{2},\gamma,\bm{Y}=\bm{y})$, for each $h\in[{K_{1}}^{*},\dots,{K_{m}}^{*}]$ is $N(\mu_{h}^{*},{\sigma_{h}^{2}}^{*})$, where $\mu_{h}^{*}={\sigma_{h}^{2}}^{*}\times\big(\sum\limits_{i:K_{i}=h}X_{i}/{\sigma_{h}^{2}}+m_{\mu}/\sigma_{z}^{2}\big)$, and ${\sigma_{h}^{2}}^{*}=\frac{{\sigma_{h}^{2}}{\sigma_{z}^{2}}}{n_{h}{\sigma_{z}^{2}}+{\sigma_{h}^{2}}}$, where $n_{h}$ is the number of times $K_{h}^{*}$ occurs in K. For $h\in\textbf{K}-[{K_{1}}^{*},\dots,{K_{m}}^{*}]$ independently simulate from $N(m_{\mu},{\sigma_{z}^{2}})$.